OpenStudy (anonymous):

solve equation (r-3) (r+6) = 0 ... someone teach me how to do these so, i don't have to keep asking questions

OpenStudy (anonymous):

r-3=0,r+6=0 r=3,r=-6

OpenStudy (anonymous):

set each term to zero (r-3)=0 (r+6)=o and solve for r r=3 r=-6

OpenStudy (anonymous):

okay i get it now lol, thank you!